8*k^2+5*k=2*k^2+4

Thanks for the help,
Chris

Answer by rapaljer(4671)

Set the equation equal to zero:
8k^2+5k=2k^2+4
8k^2 -2k^2 + 5k -4 = 2k^2 + 4 - 2k^2- 4
6k^2 + 5k - 4 = 0

Factoring this trinomial is the hard part!! Try this combination:
6k^2 + 5k - 4 = 0
(3k + 4 )(2k - 1)= 0
3k = -4 or 2k = 1
k= -4/3 or k = 1/2

R^2 at SCC